# Exploring Semi-Supervised Learning for Online Mapping

### v1

@@ Abstract
Online mapping is important for scaling autonomous driving beyond well-defined areas. Training a model to produce a local map, including lane markers, road edges, and pedestrian crossings using only onboard sensory information, traditionally requires extensive labelled data, which is difficult and costly to obtain. This paper draws inspiration from semi-supervised learning techniques in other domains, demonstrating their applicability to online mapping. Additionally, we propose a simple yet effective method to exploit inherent attributes of online mapping to further enhance performance by fusing the teacher’s pseudo-labels from multiple samples. The performance gap to using all labels is reduced from 29.6 to 3.4 mIoU on Argoverse, and from 12 to 3.4 mIoU on NuScenes utilising only 10% of the labelled data. We also demonstrate strong performance in extrapolating to new cities outside those in the training data. Specifically, for challenging nuScenes, adapting from Boston to Singapore, performance increases by 6.6 mIoU when unlabelled data from Singapore is included in training.
Figure 1: Self-training with Teacher Student set-up for Online Mapping.
I Introduction
@@
An autonomous vehicle needs to understand where it is safe to drive. This includes knowing where other road participants are and where the lane delimiters and road edges are located around the vehicle. Relying solely on pre-made offline High-Definition (HD) maps for road information comes with difficulties, such as accurate localisation within the map, and the costs associated with building and keeping the maps up to date. The latter is significant for scaling Autonomous Driving (AD) beyond limited, well-defined areas. Online mapping is an alternative and complementary approach, where onboard sensory data is used to produce a local up-to-date map depicting surrounding lane markers, road edges, and pedestrian crossings in real-time [24, 25].
@@
Recent studies have shown that end-to-end learning-based online mapping algorithms struggle to generalise to regions not present in their training set, suggesting that the available data is insufficient for learning this complex task [27, 29]. While collecting large amounts of AD data is relatively straightforward, the bottleneck lies in the effort needed to label the data. Creating the necessary labels, e.g. 3D polylines for each lane divider, is generally time-consuming and costly. Therefore, we are interested in approaches that utilize a small amount of labelled data and a vast amount of unlabelled data to increase the generalisability in the online mapping task.
@@
Other domains, such as image classification [39] and semantic segmentation [13], have addressed this problem setting through Semi-Supervised Learning (SSL) and shown remarkable capabilities. Recently the interest in SSL for object detection has also gained traction and shown promising results [32, 47]. Some SSL techniques have also been successfully adapted to the closely related domain of Bird’s Eye View (BEV) semantic segmentation. For instance, utilising a teacher-student setup [39] has demonstrated effectiveness in monocular BEV segmentation, applicable to both dynamic classes such as vehicles and pedestrians, and static classes including drivable areas, and walkways [55, 14]. However, new challenges arise when multiple cameras are used as input, as creating 3D-consistent augmentations in this context proves difficult. For instance, the conjoint-rotation technique introduced by Zhu et al. [55] is incompatible with our problem setup. Concurrent work has shown that randomly disabling entire camera feeds and applying dropout on the BEV features as strong augmentation for the student are effective when surround-view images are used as input [21].
In this work, we aim to identify which key SSL approaches excel in the online mapping domain and explore how to best leverage the unique characteristics of online mapping. For instance, the static nature of objects in online mapping allows for leveraging temporal consistency to enhance performance, which is not possible in the more general problem formulation of BEV semantic segmentation. In this spirit, we propose and investigate a label-efficient SSL method, illustrated in Fig. 1, that uses sequential data to enable the teacher to fuse information across multiple samples and thus produce more reliable pseudo-labels. This approach effectively boosts performance in online mapping leveraging less expensive unlabelled data.
@@
This paper makes four key contributions: First, we conduct a comprehensive analysis of prevalent SSL techniques for online mapping, demonstrating which existing methods can significantly enhance performance. Second, we propose using the static-world assumption to improve pseudo-label accuracy by enabling the teacher model to integrate information across multiple samples. Third, we combine SSL components into a robust method that maintains high performance levels across varying volumes of labelled data. Lastly, we demonstrate this method’s additional benefit of strong performance in domain adaptation when trained in one group of cities and evaluated in another.
II Related Work
@@
This work explores which semi-supervised learning, and domain adaptation, approaches excel when applied to the online mapping setting. Thus, we describe these fields in the following section.
II-A Semi-Supervised Learning
@@
In SSL, we combine costly labelled data with inexpensive unlabelled data to train a model. There are two main approaches to this problem, single-stage training where automatic ground truth labels are generated during training, i.e. using predictions from a single model or an ensemble of models as pseudo-labels [37, 51, 23], or a two-stage process starting with self-supervised pre-training followed by finetuning on the labelled data [8, 19, 5]. This paper concentrates on the single-stage joint training approach, exploring critical design decisions and their impact on the performance of semi-supervised online mapping.
@@
SSL has a rich history in image classification, with the teacher-student framework [39] being particularly influential. This framework involves a student network supervised by predictions from a teacher network. The teacher network shares the same architecture as the student but updates its weights as an exponential moving average (EMA) of the student’s weights. ReMixMatch [1] further introduced the idea of feeding the teacher model weakly augmented inputs, while the student model receives strongly augmented inputs. Many works have highlighted the importance of data augmentation in SSL [54, 13, 41] and these augmentations may differ from those suitable for fully supervised settings [54]. Common augmentation strategies for SSL include RandAugment [10], CutOut [11], CutMix [49], and feature dropout [48].
@@
Our setup closely resembles the concurrent work of PCT [21], which proposes SSL on multi-view images using a teacher-student framework inspired by [48]. They introduce a camera dropout augmentation technique, where one or more camera images are omitted for the student model, and leverage BEV feature dropout. Additionally, they use pre-trained models to generate PV semantic segmentation and depth pseudo-labels. However, in this paper, we focus on core SSL methods without incorporating additional pre-trained networks for pseudo-labeling.
II-B Domain Adaptation
Unsupervised domain adaptation (UDA) is a field closely related to SSL, where the unlabelled data is drawn from a target domain that is slightly shifted from the source domain (labelled data). The goal of UDA is to achieve strong performance in the target domain and is of great importance to AD as we want to generalise to new geographical regions. UDA incorporates many techniques commonly used in SSL, with self-training being the most prominent [56]. There are many works on UDA for semantic segmentation with AD applications in mind. The dominant approach in these works involves augmenting data by mixing crops of labelled and unlabelled data [40, 7]. However, such augmentations are not well-suited for our setting, as online mapping requires consistency across multiple camera views.
II-C Online Mapping
@@
We are mainly interested in the effects of SSL approaches and will use a segmentation-based method as the formulation is a better match for existing SSL and UDA methods.
III SSL for Online Mapping
Our work focuses on leveraging SSL to maximize the effectiveness of unlabelled data in online mapping. In this context, we assume that we have a small set of labelled data ${\mathbf{X}_{L}=\{\mathbf{x_{i}},\mathbf{y_{i}}\}_{i=1}^{n_{l}}}$, for which ground truth map $\mathbf{y}_{i}$ is available, and a large set of unlabelled images $\mathbf{X}_{U}=\{\mathbf{x_{i}}\}_{i=1}^{n_{u}}$ without ground truth annotations. The latter is assumed to be collected in areas separate from the labelled set, but using the same cameras. Further, the relative pose of the vehicle in the sequences is assumed to be known.
Using the labelled data, the student is trained in an ordinary supervised manner by minimizing a supervised loss ${\ell_{\text{sup}}(\mathbf{x_{i}},\mathbf{y_{i}})}$, e.g., cross-entropy, DICE [38], or Focal Loss [28]. Effectively using the unlabelled data is more challenging as we do not have access to a ground truth map for this area, and several design choices must be made. To understand how to apply SSL methods in the context of online mapping, we explore the impact of these choices using a well-established online mapping architecture. Figure 1 shows an overview of the SSL training scheme, and in this section, we provide the details of the online mapping architecture, the prevalent SSL techniques we employed, and our proposed simple multi-step teacher fusion approaches.
III-A Online mapping architecture
As we are mainly interested in the effects of SSL approaches, we choose to use a typical map segmentation architecture [24, 52, 34] such as Fig. 1 illustrates. A set of multi-view images, $\mathbf{x}$, is individually encoded by a neural network ${\mathbf{z}_{\text{pv}}=f^{\text{enc}}_{\theta}(\mathbf{x})}$ and lifted to BEV features as ${\mathbf{z}_{\text{bev}}=f^{\text{lift}}_{\theta}(\mathbf{\mathbf{z}_{\text{pv}}})}$. Then, another neural network decodes the BEV features ${\tilde{\mathbf{z}}_{\text{bev}}=f^{\text{dec}}_{\theta}(\mathbf{z}_{\text{bev}})}$. The final class probabilities are predicted through a classification head ${\mathbf{p}_{\text{cls}}=f^{\text{cls}}_{\theta}(\tilde{\mathbf{z}}_{\text{bev}})}$, where ${\mathbf{p}_{\text{cls}}}$ comprise the independent probability of each map grid cell containing a lane delimiter, road edge, or pedestrian crossing. As such, the online map is treated as a multi-label semantic segmentation problem. The complete architecture can be described as ${\mathbf{p}_{\text{cls}}=F_{\theta}(\mathbf{x})}$.
III-B Semi-supervised training scheme
@@
We investigate two common forms of SSL schemes, pseudo-labelling and consistency regularization through feature similarity. The first follows the approach in [1, 37, 39, 56], where the student is trained using pseudo-labels (pseudo-maps for our setting) predicted by the teacher. The teacher network receives a weakly augmented version of the input images $\mathbf{x}$ (and possibly a set of additional surrounding data samples, Sec. III-F), for which it predicts a pseudo-map ${\hat{\mathbf{y}}=\varphi(F_{\theta_{T}}(\mathbf{x}))}$. Here, $\varphi(\cdot)$ is used to decide/design suitable pseudo-labels based on the predicted class probabilities, e.g., thresholding and/or sharpening (see Sec. III-D). The student receives a strongly augmented version of the same sample $\tilde{\mathbf{x}}$ (see Sec. III-C) and is trained to make consistent class predictions with the teacher model through a pseudo-label loss ${\ell_{\text{cls}}(F_{\theta_{S}}(\tilde{\mathbf{x}}),\varphi(F_{\theta_{T}}(\mathbf{x})))}$.
The feature similarity supervision on the other hand enforces consistency between intermediate features for the teacher and student models. More formally, we train the student network to minimize a feature similarity loss $\ell_{\text{feat}}(\mathbf{z}_{\theta_{T}}(\mathbf{x}),\mathbf{z}_{\theta_{S}}(\tilde{\mathbf{x}}))$, where $\mathbf{z}_{\theta_{T}}$ and $\mathbf{z}_{\theta_{S}}$ are intermediate features on the same level, e.g. $\mathbf{z}_{\text{bev}}$ or $\tilde{\mathbf{z}}_{\text{bev}}$, using teacher and student models, respectively. In contrast to the pseudo-label training, feature similarity is insensitive to errors in teacher predictions as we avoid any label decisions. We only assume that the representation of the same scene should be consistent in both models. The details of this loss are discussed in Sec. III-E.
@@
III-C Augmentations
The strong augmentation techniques used for the input fed to the student are crucial [54]. Thus, we investigate various data augmentation techniques that do not change the geometry of the scene, as discussed in Sec. II-A. The augmentation techniques investigated in this work include photometric augmentation (random adjustments to brightness, contrast, color conversion, saturation, hue, and channel swaps), camera drop augmentation [21], CutOut [11], and BEV feature dropout inspired by [48].
III-D Sharpening and thresholding
@@
In SSL, predictions from the teacher are often refined through thresholding and sharpening before being used as pseudo-labels for the student. Thresholding involves keeping only predictions above some confidence level, whereas sharpening scales the predicted distribution to increase the confidence of the pseudo-label. Although there are many variations of thresholding and sharpening, we focus on a few key approaches. For thresholding, we follow FixMatch [37], i.e., keeping only predictions exceeding a fixed confidence threshold. For sharpening, we explore two techniques: (1) hard pseudo-labeling where the teacher predictions are reduced to its argmax [37, 22], and (2) temperature sharpening, where the teacher’s predicted logits are rescaled through division by a temperature parameter (denoted $T$ in [17, 2]) before applying sigmoid to get a soft pseudo-label.
III-E Feature Similarity
In addition to pseudo-labelling, we can also enforce feature similarity at the BEV feature level, i.e. adding the term $\omega_{\text{feat}}\ell_{\text{feat}}(\mathbf{z}_{\theta_{T}}(\mathbf{x}),\mathbf{z}_{\theta_{S}}(\tilde{\mathbf{x}}))$ to the total loss. This approach requires the student to maintain consistent BEV features with the teacher, which have been shown to improve performance [42]. While DoubleMatch used cosine similarity as the loss function, previous work on BEV segmentation employed Mean Squared Error (MSE) [55, 14]. We are therefore interested in exploring the impact of the chosen loss function on performance. Another design choice involves determining the optimal stage to harmonize the features between the teacher and the student. We identify two potential levels of interest: immediately following the lifting method, $\mathbf{z}_{\text{bev}}$, and just before decoding the class, $\tilde{\mathbf{z}}_{\text{bev}}$. Incorporating the feature similarity loss can potentially destabilize the training process as overemphasizing this loss may lead to a trivial solution.
III-F Teacher Fusion
Unlike BEV segmentation, which includes dynamic classes, all classes in online mapping are stationary. Since sensor data is collected in consecutive sequences with known ego vehicle movement, we can improve the teacher’s pseudo-labels by fusing predictions from multiple nearby samples. As Fig. 1 illustrates, predictions from past and future samples can be inferred and transformed into the current frame. By selecting only the most confident teacher predictions from any of the samples we allow the teacher to have a broader view of the nearby map and make more accurate pseudo-labels. This is especially powerful as predictions far from the ego vehicle tend to be less accurate [12]. As an alternative, we can fuse the BEV features across multiple samples by taking the mean of overlapping processed BEV features $\tilde{\mathbf{z}}_{\text{bev}}$.

### v2

@@ Abstract
The ability to generate online maps using only onboard sensory information is crucial for enabling autonomous driving beyond well-mapped areas. Training models for this task – predicting lane markers, road edges, and pedestrian crossings – traditionally require extensive labelled data, which is expensive and labour-intensive to obtain. While semi-supervised learning (SSL) has shown promise in other domains, its potential for online mapping remains largely underexplored. In this work, we bridge this gap by demonstrating the effectiveness of SSL methods for online mapping. Furthermore, we introduce a simple yet effective method leveraging the inherent properties of online mapping by fusing the teacher’s pseudo-labels from multiple samples, enhancing the reliability of self-supervised training. If 10% of the data has labels, our method to leverage unlabelled data achieves a 3.5x performance boost compared to only using the labelled data. This narrows the gap to a fully supervised model, using all labels, to just 3.5 mIoU. We also show strong generalization to unseen cities. Specifically, in Argoverse 2, when adapting to Pittsburgh, incorporating purely unlabelled target-domain data reduces the performance gap from 5 to 0.5 mIoU. These results highlight the potential of SSL as a powerful tool for solving the online mapping problem, significantly reducing reliance on labelled data.
Figure 1: Semi-supervised learning framework for online mapping with temporal pseudo-label fusion. Our approach adapts the Teacher-Student paradigm to the online mapping domain while introducing a novel temporal fusion mechanism that aggregates pseudo-labels across multiple frames. We enable label-efficient learning from largely unlabelled datasets and strong generalization to unseen cities.
1 Introduction
@@
Safe autonomous navigation requires accurately perceiving both dynamic traffic participants and static road infrastructure. Relying solely on pre-made offline High-Definition (HD) maps for the latter comes with difficulties, such as accurate localisation within the map, and the costs associated with building and keeping the maps up to date. The latter is significant for scaling Autonomous Driving (AD) beyond limited, well-mapped areas. Online mapping is an alternative and complementary approach, where onboard sensory data is used to produce a local up-to-date map of surrounding lane markers, road edges, and pedestrian crossings in real-time [25, 26].
@@
Recent studies have shown that learning-based online mapping algorithms struggle to generalise to regions outside their training set, suggesting available data is insufficient for learning this complex task [28, 30, 40]. While collecting large amounts of AD data is relatively straightforward, the bottleneck lies in the effort labelling the data. Creating the necessary labels, e.g. 3D polylines for each lane divider, is generally time-consuming and costly. Therefore, we are interested in approaches utilizing a small amount of labelled data and a vast amount of unlabelled data to increase the generalisability in the online mapping task.
@@
Other domains, such as image classification [43] and semantic segmentation [14], have addressed this problem setting through Semi-Supervised Learning (SSL) and shown strong capabilities. Recently, the interest in SSL for object detection has also gained traction and shown promising results [33, 51]. Some SSL techniques have been successfully adapted to the closely related domain of Bird’s Eye View (BEV) semantic segmentation of the combination of dynamic classes such as vehicles and pedestrians, and static classes including vegetation and drivable areas. However, the focus has been on monocular BEV segmentation where the teacher-student setup [43] has shown effectiveness [60, 15]. New challenges arise when multiple cameras are used as input, as creating 3D-consistent augmentations in this context proves difficult. For instance, the conjoint-rotation technique [60] is incompatible with our problem setup. Concurrent work [22] shows that randomly disabling camera feeds and applying BEV feature dropout as student augmentation techniques effectively enhance performance in a multi-camera setup. Despite these advances, there remains a gap in understanding how SSL components specifically interact in the context of online mapping, where the focus is exclusively on static classes rather than the mixed dynamic-static scenario addressed in prior work.
In this work, we identify key SSL approaches that excel in online mapping, both for learning with limited labels and for adapting models to new cities. As online mapping focuses solely on static classes, it allows us to exploit temporal consistency across frames. Building on this insight, we propose a label-efficient SSL method, illustrated in Fig. 1, that adapts the Teacher-Student paradigm to online mapping while introducing a novel temporal fusion mechanism. This mechanism aggregates pseudo-labels across multiple frames, improving their reliability and reducing noise. By leveraging large-scale unlabelled datasets, our approach significantly reduces dependence on labelled data while enabling strong generalization to unseen cities.
@@
To summarize, we make four key contributions: 1) a comprehensive analysis of prevalent SSL techniques for online mapping, demonstrating which existing methods can significantly enhance performance. 2) combine SSL components into a robust method that maintains high performance levels across varying volumes of labelled data. 3) propose that the teacher model integrates information across multiple samples to improve pseudo-label accuracy. 4) demonstrate strong performance in domain adaptation when trained in one group of cities and evaluated in another.
2 Related Work
@@
We explore semi-supervised learning and domain adaptation techniques specifically applied to online mapping. This section examines relevant prior work in these fields.
2.1 Semi-Supervised Learning
@@
In SSL, costly labelled data is combined with inexpensive unlabelled data to train a model. There are two main approaches to this problem: single-stage training, where automatic ground truth labels are generated during training, i.e. using predictions from a single model or an ensemble of models as pseudo-labels [41, 55, 24], or a two-stage process starting with self-supervised pre-training followed by finetuning on the labelled data [9, 21, 5, 34]. This paper concentrates on the single-stage joint training approach, exploring critical design decisions and their impact on the performance of semi-supervised online mapping.
@@
SSL has a rich history in image classification, with the teacher-student framework [43] being particularly influential. This framework involves a student network supervised by predictions from a teacher network. The teacher network shares the same architecture as the student but updates its weights as an exponential moving average (EMA) of the student’s weights. ReMixMatch [1] further feeds the teacher weakly augmented inputs, while the student receives strongly augmented inputs. Data augmentation is crucial in SSL [58, 14, 45] and these may differ from those suitable for fully supervised settings [58]. Commonly used augmentation strategies include RandAugment [11], CutOut [12], CutMix [53], and feature dropout [52].
@@
Our setup closely resembles the concurrent work of PCT [22], which proposes SSL on multi-view images using a teacher-student framework inspired by [52]. They introduce camera dropout augmentation, where one or more camera images are omitted for the student model, and leverage BEV feature dropout. Additionally, they use pre-trained models to generate PV semantic segmentation and depth pseudo-labels. In contrast, our work focuses on SSL methods that do not rely on additional pre-trained networks for pseudo-labeling. Another key distinction is that, unlike BEV segmentation, which includes dynamic object classes, online mapping is concerned solely with stationary elements.
2.2 Domain Adaptation
Unsupervised domain adaptation (UDA) is closely related to SSL, and throughout this paper, we will use the terms interchangeably. In UDA, the unlabelled data comes from a target domain that differs from the labelled source domain. The goal of UDA is to achieve strong performance in unseen environments, which is particularly critical for online mapping, where models must perform well in cities they have never been trained on – a problem often overlooked in prior work [28]. Similar to SSL, many UDA techniques rely on self-training with a teacher-student setup and have also demonstrated strong results across various UDA settings [61]. A dominant approach for perspective view semantic segmentation in the image plane for AD involves augmenting data by mixing crops of labelled and unlabelled samples [44, 7]. However, such augmentations are poorly suited for online mapping, as they fail to maintain consistency across multiple camera views, a key requirement in our setting. In this work, we focus on self-training for UDA in online mapping, emphasizing its importance in adapting to unseen cities – a fundamental challenge for real-world deployment.
2.3 Online Mapping
@@
Previous works have highlighted a performance drop as the domain gap between training and evaluation data increases. Particularly in the context of geographical distribution shifts [28, 40], and driving scenario shifts [30]. Our focus is on the effects of SSL approaches, and for this work, we choose a segmentation-based method, as its formulation aligns more naturally with existing SSL and UDA techniques.
3 SSL for Online Mapping
Our work focuses on leveraging SSL to maximize the effectiveness of unlabelled data in online mapping. In this context, we assume that we have a small set of labelled data ${\mathbf{X}_{L}=\{\mathbf{x_{i}},\mathbf{y_{i}}\}_{i=1}^{n_{l}}}$, for which ground truth map $\mathbf{y}_{i}$ is available, and a large set of unlabelled images $\mathbf{X}_{U}=\{\mathbf{x_{i}}\}_{i=1}^{n_{u}}$ without ground truth annotations. The latter is assumed to be collected in areas separate from the labelled set, but using the same camera setup. Further, the relative pose of the vehicle in the sequences is assumed to be known.
Using the labelled data, the student is trained in an ordinary supervised manner by minimizing a supervised loss ${\ell_{\text{sup}}(\mathbf{x_{i}},\mathbf{y_{i}})}$, e.g., cross-entropy, DICE [42], or Focal Loss [29]. Effectively using the unlabelled data is more challenging as we do not have access to a ground truth map for this area, and several design choices must be made. To understand how to apply SSL methods in the context of online mapping, we explore the impact of these choices using a well-established segmentation-based online mapping architecture. Figure 1 shows an overview of the SSL training scheme, and in this section, we provide the details of the online mapping architecture, the prevalent SSL techniques we employed, and our proposed multi-step teacher fusion approaches.
3.1 Online mapping architecture
As we are mainly interested in the effects of SSL approaches, we choose to use a typical map segmentation architecture [25, 56, 37] such as Fig. 1 illustrates. The multi-view images, $\mathbf{x}$, are encoded by a neural network ${\mathbf{z}_{\text{pv}}=f^{\text{enc}}_{\theta}(\mathbf{x})}$ and lifted to BEV features as ${\mathbf{z}_{\text{bev}}=f^{\text{lift}}_{\theta}(\mathbf{\mathbf{z}_{\text{pv}}})}$. Then, another neural network decodes the BEV features ${\tilde{\mathbf{z}}_{\text{bev}}=f^{\text{dec}}_{\theta}(\mathbf{z}_{\text{bev}})}$. Final class probabilities are predicted through a classification head ${\mathbf{p}_{\text{cls}}=f^{\text{cls}}_{\theta}(\tilde{\mathbf{z}}_{\text{bev}})}$, where ${\mathbf{p}_{\text{cls}}}$ comprise the independent probability of each map grid cell containing a lane delimiter, road edge, or pedestrian crossing. As such, the online map is treated as a multi-label semantic segmentation problem. The complete architecture can be described as ${\mathbf{p}_{\text{cls}}=F_{\theta}(\mathbf{x})}$.
3.2 Semi-supervised training scheme
@@
We investigate two common forms of SSL schemes, pseudo-labelling and consistency regularization through feature similarity. The first follows the approach in [1, 41, 43, 61], where the student is trained using pseudo-labels (pseudo-maps for our setting) predicted by the teacher. The teacher network receives a weakly augmented version of the input images $\mathbf{x}$ (and possibly a set of additional surrounding data samples, Sec. 3.6), for which it predicts a pseudo-map ${\hat{\mathbf{y}}=\varphi(F_{\theta_{T}}(\mathbf{x}))}$. Here, $\varphi(\cdot)$ is used to decide/design suitable pseudo-labels based on the predicted class probabilities, e.g., thresholding and/or sharpening (see Sec. 3.4). The student receives a strongly augmented version of the same sample $\tilde{\mathbf{x}}$ (see Sec. 3.3) and is trained to make consistent class predictions with the teacher model through a pseudo-label loss ${\ell_{\text{cls}}(F_{\theta_{S}}(\tilde{\mathbf{x}}),\varphi(F_{\theta_{T}}(\mathbf{x})))}$ similar to $\ell_{\text{sup}}$.
The feature similarity supervision, on the other hand, enforces consistency between intermediate features for the teacher and student models. More formally, we train the student network to minimize a feature similarity loss, using e.g. Mean Squared Error or Cosine Similarity, $\ell_{\text{feat}}(\mathbf{z}_{\theta_{T}}(\mathbf{x}),\mathbf{z}_{\theta_{S}}(\tilde{\mathbf{x}}))$. Here, $\mathbf{z}_{\theta_{T}}$ and $\mathbf{z}_{\theta_{S}}$ are intermediate features on the same level, e.g. $\mathbf{z}_{\text{bev}}$ or $\tilde{\mathbf{z}}_{\text{bev}}$, using teacher and student models, respectively. In contrast to the pseudo-label training, feature similarity is insensitive to errors in teacher predictions as we avoid any label decisions. Instead, we enforce the representation of the same scene to be consistent in both models and is further discussed in Sec. 3.5.
@@
3.3 Augmentations
Strong augmentation techniques applied to the student’s input are essential for effective learning [58]. Therefore, we explore various data augmentation strategies that preserve the scene’s geometry, as discussed in Sec. 2.1. Specifically, we employ photometric augmentations, CamDrop [22], CutOut [12], and BEVDrop [52]. Photometric augmentations introduce random variations in appearance by adjusting brightness, contrast, saturation, hue, and color conversion, as well as performing random channel swaps. CamDrop augmentation simulates missing camera views by completely removing the input from one or more cameras. The dropped camera’s field of view is also masked from the loss. CutOut randomly masks out rectangular regions in the input images while BEVDrop operates on the BEV feature representation randomly setting certain grid cells to zero.
3.4 Sharpening and thresholding
@@
Predictions from the teacher are often refined through thresholding and sharpening before being used as pseudo-labels for the student. Thresholding keeps only predictions above some confidence level, whereas sharpening scales the predicted distribution to alter the confidence of the pseudo-label. For thresholding, we adapt FixMatch [41], i.e., only predictions exceeding a fixed confidence threshold contribute to the loss. For sharpening, we explore two techniques: (1) hard pseudo-labeling where teacher predictions are reduced to its argmax [41, 23], and (2) temperature sharpening, where the predicted logits are rescaled through division by a temperature parameter (denoted $T$ in [18, 2]) before applying sigmoid to get a soft pseudo-label. More formally, splitting up the classification head, $\mathbf{p}_{\text{cls}}=f^{\text{cls}}_{\theta}(\mathbf{\tilde{z}}_{\text{bev}})=\sigma(f^{\text{logits}}_{\theta}(\tilde{\mathbf{z}}_{\text{bev})})$, we have the intermediate logits $\mathbf{\hat{z}}$. Sharpening using the hyperparameter $T$ is then defined as $\mathbf{\hat{z}}_{sharp}=\mathbf{\hat{z}}/T$. For $10<T<$ 1 the certainty is artificially increased, i.e. $\mathbf{\hat{z}}_{sharp}>\mathbf{\hat{z}}$. Conversely, $T>1$ decreases the pseudo-label certainty.
3.5 Feature Similarity
To further enhance consistency between the teacher and student networks, we impose feature similarity constraints at the BEV feature level. A key consideration in this approach is selecting both the loss function and the feature level at which to apply the constraint. We explore two candidate feature levels for enforcing similarity: Early BEV representation ($\mathbf{z}_{\text{bev}}$) – immediately after the lifting process, capturing the initial transformation of multi-view image features into the BEV space. Late BEV representation ($\tilde{\mathbf{z}}_{\text{bev}}$) – just before the final decoding step, where the network has refined its scene understanding. By constraining intermediate features to be consistent, we encourage the student model to align with the teacher’s representations without relying on explicit pseudo-labels. However, excessive weighting of the feature similarity loss may destabilize training, potentially leading to trivial solutions where the student merely replicates the teacher’s features instead of learning meaningful representations.
3.6 Teacher Fusion
Unlike BEV segmentation, which includes dynamic classes, all categories in online mapping are stationary. This key distinction, combined with the fact that sensor data is collected in sequential frames with known ego vehicle motion, enables novel strategies not commonly explored in other domains. We propose to leverage this by fusing pseudo-labels from the teacher across multiple nearby frames.
As illustrated in Fig. 1, predictions from both past and future frames can be transformed into the current coordinate frame using the known ego motion. By selecting only the most confident teacher predictions from any of these frames, we provide the student with a more complete and accurate map representation. This approach is particularly beneficial for improving predictions in regions far from the ego vehicle, where accuracy tends to degrade [13]. We also explore feature fusion by averaging overlapping BEV feature representations ($\tilde{\mathbf{z}}_{\text{bev}}$) across multiple frames.